OpenStudy (anonymous):

The first four terms in a quadratic sequence are shown below. 5, 15, 31, 53, ... What is the next term in the sequence?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

differences are 15-5=10 31-15=16 53-31=22 that means the next term will be 53+28 = 81

OpenStudy (anonymous):